zarkam21 (zarkam21):

Help with 3 part question

OpenStudy (oliviadaae):

Could you post the four parts please? :)

zarkam21 (zarkam21):

Actually it's 2 parts =)

OpenStudy (oliviadaae):

College Algebra I'm guessing? I'm currently in this class, and I haven't learned this yet. So sorry.

OpenStudy (danjs):

did you get anything yey?

zarkam21 (zarkam21):

Nope

OpenStudy (danjs):

They tell you the standard equation for an ellipse, center at (h,v) \[\frac{ (x-h)^2 }{ a^2 }+\frac{ (y-v)^2 }{ b^2 }=1\]

zarkam21 (zarkam21):

@3mar

OpenStudy (3mar):

Well, I am here.

OpenStudy (danjs):

Your ellipse for this problem is... compare that to the standard form, the center is (h,k) \[\large \frac{ (x+3)^2 }{ 9 }+\frac{ y^2 }{ 25 }=1\] you can make it look more the same \[\large \frac{ (x-(-3))^2 }{ 3^2 }+\frac{ (y-0)^2 }{ 5^2 }=1\]

OpenStudy (danjs):

center (h,k)=(-3 , 0), right?

zarkam21 (zarkam21):

right

OpenStudy (3mar):

@zarkam21 With DanJS, you are in good hands! Thank you for good illustration, DanJS!

zarkam21 (zarkam21):

Thank you @3mar

OpenStudy (3mar):

Don't mention that!

OpenStudy (danjs):

the axis have to do with the values of 'a' and 'b' 'b' is the y direction, and 'a' is the x direction

OpenStudy (danjs):

Center (h,v)=(-3 , 0) a = 3 b = 5